### v2

# Spatial Re-parameterization for N:M Sparsity

Abstract
This paper presents a Spatial Re-parameterization (SpRe) method for the N:M sparsity. SpRe stems from an observation regarding the restricted variety in spatial sparsity presented in N:M sparsity compared with unstructured sparsity. Particularly, N:M sparsity exhibits a fixed sparsity rate within the spatial domains due to its distinctive pattern that mandates N non-zero components among M successive weights in the input channel dimension of convolution filters. On the contrary, we observe that conventional unstructured sparsity displays a substantial divergence in sparsity across the spatial domains, which we experimentally verify to be very crucial for its robust performance retention compared with N:M sparsity. Therefore, SpRe employs the spatial-sparsity distribution of unstructured sparsity by assigning an extra branch in conjunction with the original N:M branch at training time, which allows the N:M sparse network to sustain a similar distribution of spatial sparsity with unstructured sparsity. During inference, the extra branch can be further re-parameterized into the main N:M branch, without exerting any distortion on the sparse pattern or additional computation costs. SpRe has achieved a commendable feat by matching the performance of N:M sparsity methods with state-of-the-art unstructured sparsity methods across various benchmarks. Our project is available at https://github.com/zyxxmu/SpRE.
Index Terms:
Convolutional neural networks, Network sparsity, N:M sparsity.
1 Introduction
Network sparsity has proven many successes in reducing the complexity of convolutional neural networks (CNNs) [1, 2, 3]. A sparse network can be obtained by zeroizing weights at different granularity levels, from fine to coarse. Fine-grained sparsity (unstructured sparsity) [2, 4] removes individual weights and is demonstrated to well retain performance even at high sparsity rates. Unfortunately, the deployment of fine-grained sparse networks on off-the-shelf hardware is cumbersome due to the irregularity of sparse weight matrices. On the other hand, coarse-grained sparsity (structured sparsity) [5, 6] achieves significant acceleration by eliminating entire convolution filters [7, 6] or weight blocks [8, 9], yet suffers severe performance degradation particularly at high sparsity rates.
Figure 1: A toy example of the discrepancy in spatial sparsity between N:M sparsity at 1:4 pattern and unstructured sparsity at 75% sparsity rate. (a) N:M sparsity requires N non-zero components among M consecutive weights in the input channel dimension, resulting in equal spatial sparsity. (b) Unstructured sparsity removes weights at arbitrary locations, resulting in uneven spatial sparsity.
N:M sparsity has lately surfaced as a promising direction of augmenting the trade-off between acceleration effects and performance retention [10, 11]. By stipulating at most N non-zero components within M consecutive weights across the input channel dimension, it considerably improves the performance of structured sparsity while simultaneously ensuring expeditious inference aided by the N:M sparse tensor core [12]. In recent years, various methods have been proposed to train N:M sparse networks from pre-trained weights [12, 11] or randomly-initialized weights [10, 13]. Nevertheless, the performance of N:M sparsity still lags behind unstructured sparsity, particularly at high sparsity rates such as 95% [14, 10]. We ask: what leads to the performance gap between N:M sparsity and unstructured sparsity?
In this paper, we answer this question through empirical observation of network sparsity over the spatial domain. Particularly, N:M sparsity displays a consistent sparsity rate of $1-\frac{\text{N}}{\text{M}}$ at every spatial location of convolution filters owing to the distinctive sparse pattern depicted in the input channel dimension (Figure 1a). Conversely, unstructured sparsity can exhibit a notable variation in spatial sparsity (Figure 1b), which we confirm to be ubiquitous in existing unstructured sparsity methods and crucial for their robust performance retention compared with N:M sparsity methods (Sec. 3.2). To explain this, a heterogeneous distribution of weights across various positions aids in giving precedence to the most informative visual elements in the spatial domain, consequently resulting in enhanced performance for the sparse networks. Self-evidently, N:M sparsity falters to assign adequate weights for the informative visual elements especially when the sparsity rate is high, therefore leading to more performance degradation.
Driven by this analysis, we present Spatial Re-parameterization (SpRe) as a way of matching the performance between N:M sparsity and unstructured sparsity. SpRe utilizes the spatial sparsity distribution of unstructured sparsity to allocate an extra weight branch in conjunction with the original N:M branch. This enables the sparse network unifying these two branches to maintain a sparsity distribution comparable to that of unstructured sparsity. Moreover, we constrain the newly introduced parameters to adhere to the N:M sparse distribution of the main branch in the input channel dimension. This results in an advantage for a re-parameterization after training, where the newly-added branch can be merged into the main block without impacting the output at the inference stage. Thus, SpRe introduces no additional inference burden for the original N:M sparse networks. The advantages of our proposed SpRe include:
  • Trackable. SpRE is traceable in principle, due to our innovative observation of the discrepancy in spatial sparsity between N:M sparsity and unstructured sparsity, which we prove to be essential for performance retention.
  • Scalable. SpRe is easy to use and orthogonal to enhance the effectiveness of other N:M sparsity methods, whether applied from randomly-initialized or pre-trained weights.
  • High-performance. SpRe is validated to be highly successful in boosting the performance of N:M sparsity methods across various benchmarks. Specifically, SpRe enhances the Top-1 accuracy of SR-STE [10], a leading N:M method, by 1.2% when training a 1:16 sparse ResNet-50 [15] on ImageNet [16]. Moreover, the boosted performance even surpasses state-of-the-art unstructured sparsity method GraNet [14] by 0.4% at a similar sparsity rate.
2 Related Work
2.1 Unstructured Sparsity
By removing individual weights at arbitrary positions of the network to address the heavy parameter burden, unstructured sparsity has emerged as a fervent area of research over the last decade [2, 1, 17]. Gradient [18], momentum [4], magnitude [1], and many others, are often used to identify and remove insignificant weights. Recent advancements learn to train an unstructured sparse network in a dynamic sparse manner towards performance enhancement. RigL [19] alternatively removes and revives weights based on magnitudes and gradients. Sparse Momentum [20] considers the mean momentum magnitude in each layer to redistribute weights. Besides, gradual sparsity is widely adopted to boost performance [21, 14]. Unstructured sparsity is demonstrated to well retain performance, even at a very high sparsity over 95% [14]. However, the resulting irregular sparse tensors make it gain rare speedups on general hardware [22].
2.2 N:M Sparsity
N:M sparsity preserves N out of M consecutive weights along the input channel dimension of CNNs, and achieves practical speedups thanks to the hardware innovation of N:M sparse tensor core [12, 23]. The pioneering ASP [12] goes through model pre-training, high-magnitude weight removal [1], and model fine-tuning. Pool et al. [11] proposed channel permutation to increase the performance of ASP. Sun et al. [24] proposed a layerwise fine-grained N:M sparsity to replace the common uniform version. To avoid heavy burden on model pre-training, Zhou et al. [10] proposed a sparse-refined straight-through estimator (SR-STE) to learn from scratch. More specifically, N-out-of-M weights of higher magnitudes are selected in each forward pass, whileall weights are updated during the backward phase, utilizing the STE estimator, paired with a uniquely designed sparse penalty term. LBC [13] reformulates the N:M sparsity as a combinatorial problem and learns the best combination for the sparse weights. Recent advances [25, 26] also excavate into the training efficiency of N:M sparsity through devising masks that enable N:M sparse computation in both forward and backward propagation.
2.3 Structural Re-parameterization
Structural re-parameterization mutually converts different architectures through an equivalent transformation of parameters. The representative RepVGG [27] merges kernels of smaller sizes to these of larger ones in inference. For example, 1$\times$1 kernels can be added onto the central points of the 3×3 kernels. Along this line, researchers have devised various blocks to boost the performance of a regular CNN without extra inference costs, e.g., Asymmetric Convolution Block (ACB)  [28], RepLKNet [29]. Besides, structural re-parameterization is also leveraged to guide channel pruning [30], where the original CNN is re-parameterized into two parts to respectively maintain the performance and prune convolutional filters.
In this paper, we focus on developing an N:M sparsity method that is orthogonal to the aforementioned methods for performance improvement, simply due to two advantages of our method: First, we utilize the distribution of unstructured sparsity across the spatial domain. Second, we re-parameterize a newly-added branch into the main branch of N:M block in the inference.
Figure 2: Spatial sparsity of common unstructured sparsity methods including Magnitude-based sparsity [1], RigL [19], GraNet [14]. We show spatial sparsity of 3$\times$3 kernels from different layers of ResNet-50 [15] with overall 95% sparsity, close to the sparsity level of 1:16 pattern. Experiments are performed on ImageNet-1K [16].
3 Methodology
3.1 Background
For convolutional weights $\mathbf{W}\in\mathbb{R}^{C_{o}\times C_{i}\times K\times K}$ ($C_{o}$: output channel, $C_{i}$: input channel, $K$: kernel size), the convolution operation with input features $\mathbf{X}\in\mathbb{R}^{H\times W\times C_{i}}$ ($H$: height, $W$: width) is generally formulated as:
$\mathbf{Y}=BN(\mathbf{W}\circledast\mathbf{X}),$ (1)
[…]
$\mathbf{Y}=BN((\mathbf{B}\odot\mathbf{W})\circledast\mathbf{X}),$ (2)
where $\odot$ denotes the element-wise multiplication. It is easy to know that, $\mathbf{B}_{p,q,u,v}$ = 00 removes $\mathbf{W}_{p,q,u,v}$ and $\mathbf{B}_{p,q,u,v}$ = $1$ preserves $\mathbf{W}_{p,q,u,v}$.
For unstructured sparsity, the zero entries are irregular in the positions of $\mathbf{B}$. Instead, N:M sparsity stipulates N non-zero entries for every M consecutive weights along the input channel dimension. Therefore, $\mathbf{B}$ is restricted to satisfy:
${\|\mathbf{B}_{p,\lfloor q/\text{M}\rfloor\cdot\text{M}:\lfloor q/\text{M}\rfloor\cdot\text{M}+\text{M},u,v}\|}_{0}=\text{N},$ (3)
where $p,q,u,v$ enumerates $C_{o},C_{i},K,K$, respectively.
Besides, for ease of the following representation, we define Spatial Sparsity (SS) to measure the weight sparsity in the spatial domain. The spatial sparsity at position ($u$, $v$) is calculated as:
$\text{SS}(\mathbf{B},u,v)=1-\frac{1}{C_{o}\cdot C_{i}}\sum_{p=1}^{C_{o}}\sum_{q=1}^{C_{i}}\mathbf{B}_{p,q,u,v}.$ (4)
3.2 Discrepancy in Spatial Sparsity
Compared to unstructured sparsity, N:M sparsity achieves practical speedups with the support of the N:M sparse tensor core [12, 23], yet it suffers more performance degradation, particularly at high sparsity rates [10, 14]. In this section, we demonstrate that such a performance gap stems from a discrepancy in the spatial sparsity between N:M sparsity and unstructured sparsity. Simply due to the constraint of Eq. (3), N:M sparsity possesses consistent spatial sparsity across different positions as:
$\text{SS}(\mathbf{B},u,v)=1-\frac{\text{N}}{\text{M}}.$ (5)
In other words, N:M sparsity equally allocates weights across the spatial domain of convolution operation. On the other hand, such balanced spatial sparsity is not imperative for unstructured sparsity, given that no sparsity restriction is imposed across the channel dimension. Indeed, we observe that unstructured sparsity exhibits significant variability in terms of spatial sparsity, as illustrated by Figure 2. This variability in spatial sparsity remains consistent across different layers and network types, irrespective of the specific unstructured sparsity technique employed [1, 19, 32].
Upon closer examination of Figure 2, it can be inferred that the majority of unstructured weights even maintain a similar distribution spatial sparsity, more interestingly resembling a cross-like configuration in the shallow layers. This intriguing phenomenon warrants further exploration and investigation within the domain of unstructured sparsity. We do not delve deeply into this matter at present, but it is unequivocal that unstructured sparsity methods [1, 19, 32] allocate more weights to certain fixed visual points simultaneously.
Figure 3: Performance comparison for pruning ResNet-32 [15] on CIFAR-10 [33] and ResNet-50 [15] on ImageNet-1K [16]. The involved methods include the dense version of ResNet-50 (Baseline), SR-STE [10] (N:M sparsity), unstructured sparse GraNet [14], and GraNet with two types of sparsity constraints: (1) The same spatial sparsity along the spatial dimension (GraNet), (2) Equal mask flexibility with (1) but no spatial sparsity constraint (GraNet). Performance drops when unstructured sparse method is confined to having the same spatial sparsity, even if the mask flexibility remains the same.
CNNs analyze visual imagery by sliding filters along input features to empower each weight to interact with different visual regions. We conjecture that unstructured sparsity methods implicitly acquire weights in the spatial domain that can better tell importance of input visual regions, such that the sparse network learns to focus more on crucial visual areas. To verify our hypothesis, we destroy the variable spatial sparsity of the state-of-the-art unstructured sparsity method GraNet [14] by constraining the sparsity at any location ($u$, $v$) to be the same. In particular, given a target sparsity rate $P$, we restrict the binary mask $\mathbf{B}$ to satisfy:
${\|\mathbf{B}_{:,:,u,v}\|}_{0}\leq\lfloor{P\cdot\|\mathbf{W}_{:,:,u,v}\|}_{0}\rfloor.$ (6)
The above constraint leads to the same spatial sparsity at any location, as well as limiting the flexibility of pruning masks, which refers to the solution space for network sparsity and is a key factor for performance preservation [34, 35]. For a more rigorous comparison, we provide another sparsity constraint by firstly reshaping $\mathbf{W}\in\mathbb{R}^{C_{o}\times C_{i}\times K\times K}$ into $\mathbf{W}^{\prime}\in\mathbb{R}^{K\times K\times C_{o}\times C_{i}}$, then the binary mask $\mathbf{B}$ is restricted as:
${\|\mathbf{B}_{p,q,:,:}\|}_{0}\leq\lfloor{P\cdot\|\mathbf{W}^{\prime}_{p,q,:,:}\|}_{0}\rfloor.$ (7)
Eq. (6) and Eq. (7) have the same mask flexibility, except that the former enforces sparsity in the spatial domain. Then, we perform experimental comparison in Figure 3. Compared to the vanilla unstructured method [1], limited sparsity flexibility incurs performance degradation, even close to N:M method SR-STE [10]. Furthermore, requirement of spatial sparsity by Eq. (6) is not a patch on that by Eq. (7). To explain, when weights are scarce at elevated sparsity rates, it becomes more imperative to direct the constrained weights towards more critical visual points in order to guarantee performance retention. Conversely, if spatial sparsity is maintained at an equilibrium level, these vital visual points cannot be efficiently processed, thereby leading to more performance degradation. Therefore, we can ascertain the core role of spatial sparsity variability in improving the performance of N:M methods.
Figure 4: Framework of the proposed SpRe. (a) An extra branch is built upon the spatial sparsity arising from unstructured sparse weights and the N:M sparse mask (Eq. (8)). (b) The extra branch is trained in conjunction with the main N:M branch to reimburse the spatial sparsity. After training, a re-parameterization is performed to merge these two branches, without altering the output.
3.3 Spatial Re-parameterization
We propose Spatial Re-parameterization (SpRe), as depicted in Figure 4, as a way of attaining comparable performance in N:M sparsity to that of unstructured sparsity. To address the issue of limited spatial sparsity variation in N:M, SpRe introduces an extra branch $\mathbf{B}^{S}\odot\mathbf{W}^{S}$ in conjunction with the original N:M sparse weights $\mathbf{B}\odot\mathbf{W}$. Specifically, we first obtain mask $\mathbf{B}^{U}$ of unstructured sparsity with a sparsity rate $P=1-\frac{\text{N}}{\text{M}}$. Here we utilize the classic magnitude-based pruning [1], while other metrics for unstructured sparsity [32, 18, 36] can also adapt well. Given that N:M weights $\mathbf{B}\odot\mathbf{W}$ maintain a constant spatial sparsity of rate $1-\frac{\text{N}}{\text{M}}$, we allocate less parameter burden in the extra unstructured branch as:
$\mathbf{B}^{S}_{:,:,u,v}=\left\{\begin{array}[]{ll}\mathbf{B}_{:,:,u,v},\;\textrm{if}\;\text{SS}(\mathbf{B}^{U},u,v)<1-\frac{\text{N}}{\text{M}},\\ \mathbf{0},\;\textrm{otherwise,}\end{array}\right.$ (8)
in which $u,v=1,...,K$. Then, the regular forward propagation changes from Eq. (2) to:
$\mathbf{Y}=BN((\mathbf{B}\odot\mathbf{W})\circledast\mathbf{X})+BN((\mathbf{B}^{S}\odot\mathbf{W}^{S})\circledast\mathbf{X}).$ (9)
Consequently, the total spatial sparsity, unifying these two branches, is equivalently compensated to a correspondingly leveled unstructured sparsity during the training phase. Such a practice allows the sparse network to emphasize crucial visual regions, thereby yielding enhanced performance Next, we demonstrate that the extra branch can be re-parameterized into the main branch without incurring additional parameters or computation burden at the test stage. Particularly, in each branch, the BN layer and its preceding convolution layer can be converted into a single convolution kernel, which is a very widely-used technique in the deployment of CNN networks [27]. Denoting the converted weight in the main and extra branches as $\mathbf{B}\odot\bar{\mathbf{W}}$ and $\mathbf{B}^{S}\odot\bar{\mathbf{W}}^{S}$, the final conv kernel $\mathbf{W}^{*}$ is derived by point-wise addition of the converted weights as:
$\mathbf{W}^{*}=\mathbf{B}\odot\bar{\mathbf{W}}+\mathbf{B}^{S}\odot\bar{\mathbf{W}}^{S}.$ (10)
Given that the mask of the extra branch is a subset of the main branch ($\mathbf{B}^{S}$ is a subset of $\mathbf{B}$), as per Eq. (8), the merged weights are still in the N:M form for inference. Eventually, the forward propagation at the test stage is conducted with N:M sparse weights as:
$\mathbf{Y}=\mathbf{W}^{*}\circledast\mathbf{X}.$ (11)
Self-evidently, neither interference on the N:M pattern nor extra inference burden is introduced by SpRe.
Implementation of SpRe. SpRe is orthogonal to existing N:M sparsity techniques and can be seamlessly combined to enhance their lack of spatial sparsity variety. For techniques that attain N:M sparsity after pre-training the dense weights [1, 11], SpRe can be directly applied to build the re-parameterization branch on the basis of pre-trained weights. After fine-tuning, we employ the re-parameterization to derive the N:M sparse weights for inference. For methods that excavate N:M sparsity from scratch in a sparse training manner [10, 13], SpRe can also dynamically adapt the binary mask $\mathbf{B}^{S}$ of the extra branch by looking at the updated weights during training. This ensures that the extra branch can simultaneously sustain proper spatial sparsity distribution and re-parameterization capability. As a result, SpRe can be readily integrated into existing N:M sparsity methods to gain a consistent enhancement in their performance, which is experimentally substantiated in the subsequent section. We outline the workflow for utilizing SpRe to enhance N:M sparsity methods in Alg. 1.
Require :  Network weights $\mathbf{W}$, total training epoch $\tau$.
1 $\mathbf{W}^{S}$leftarrow randomly initialization; for $k$leftarrow 1, 2, $\dots$, $\tau$  do
2       for each training step $t$  do
             Obtain N:M sparse mask $B$; // Depending on N:M method choosen
3             Obtain spatial mask $\mathbf{B}^{S}$ via Eq. (8); Forward and backward propagation via Eq. (9); Update $\mathbf{W},\mathbf{W}^{S}$ using SGD optimizer;
4       end for
5
6 end for
7Obtain the merged weights $\mathbf{W}^{*}$ via Eq. (10);
Output: Trained N:M sparse weights $\mathbf{W}^{*}$.
Algorithm 1 Spatial Re-parameterization for N:M Sparsity.
TABLE I: Results for sparsifying ResNet-32 at different N:M patterns on CIFAR-10 .
Model Method N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc
ResNet-32 Baseline - 95.2 - 95.2 - 95.2 - 95.2
ResNet-32 ASP [12] 2:4 95.0 1:4 94.3 1:8 93.5 1:16 92.6
ResNet-32 w. SpRe 2:4 95.1(+0.1) 1:4 94.7(+0.4) 1:8 94.3(+0.8) 1:16 93.7(+1.1)
ResNet-32 SR-STE [10] 2:4 94.9 1:4 94.3 1:8 93.4 1:16 92.6
ResNet-32 w. SpRe 2:4 95.1(+0.2) 1:4 94.6(+0.3) 1:8 94.3(+0.9) 1:16 93.9(+1.3)
TABLE II: Results for sparsifying ResNet-18 at different N:M patterns on ImagNet-1K.
Model Method N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc
ResNet-18 Baseline - 70.9 - 70.9 - 70.9 - 70.9
ResNet-18 ASP [12] 2:4 70.6 1:4 69.1 1:8 67.6 1:16 65.0
ResNet-18 w. SpRe 2:4 70.8(+0.2) 1:4 69.6(+0.5) 1:8 68.2(+0.6) 1:16 65.5(+0.4)
ResNet-18 SR-STE [10] 2:4 71.2 1:4 69.2 1:8 67.2 1:16 64.9
ResNet-18 w. SpRe 2:4 71.3(+0.1) 1:4 69.9(+0.7) 1:8 68.0(+0.8) 1:16 65.7(+0.8)
TABLE III: Results for sparsifying ResNet-50 at different N:M patterns on ImagNet-1K.
Model Method N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc
ResNet-50 Baseline - 77.2 - 77.2 - 77.2 - 77.2
ResNet-50 ASP [12] 2:4 77.4 1:4 76.5 1:8 75.6 1:16 71.5
ResNet-50 w. SpRe 2:4 77.7(+0.3) 1:4 76.8(+0.3) 1:8 76.2(+0.6) 1:16 72.3(+0.8)
ResNet-50 SR-STE [10] 2:4 77.0 1:4 75.3 1:8 73.8 1:16 71.5
ResNet-50 w. SpRe 2:4 77.2(+0.2) 1:4 76.1(+0.8) 1:8 74.7(+0.9) 1:16 72.7(+1.2)
ResNet-50 LBC [13] 2:4 77.2 1:4 75.9 1:8 74.0 1:16 71.8
ResNet-50 w. SpRe 2:4 77.3(+0.1) 1:4 76.4(+0.5) 1:8 74.8(+0.8) 1:16 72.9(+1.1)
4 Experiments
4.1 Settings
Our experiments include three computer vision tasks encompassing image classification on the CIFAR-10 [33] and ImageNet-1K datasets [16], object detection and instance segmentation on the COCO benchmark [37]. The specific experimental settings are respectively expounded as follows. We validate the efficacy of SpRe in elevating the performance of prominent N:M sparsity methods, including ASP [12], SR-STE [10], and LBC [13]. When SpRe is particularly applied to a N:M method, the extra block is initialized and trained with the same configuration as the main block. Our experiments cover a wide range of N:M patterns including 2:4, 1:4, 1:8, and 1:16. For image classification, we sparsify ResNet-32 [15] on CIFAR-10 dataset, and ResNet-18 [15], ResNet-50 [15], MobileNet-V1 [38], MobileNet-V2 [39] on ImageNet-1K dataset. Besides, we exploit the efficacy of SpRe to aid SR-STE [10] in training 2:4, 1:4, 1:8, and 1:16 sparse Faster RCNN [40] for object detection and Mask RCNN [41] for instance segmentation, utilizing ResNet-50 as the backbone. Following previous works, we train all networks for 300 epochs on CIFAR-10, with a weight decay of 0.005. On ImageNet, 120 epochs are given for ResNet and 300 epochs for DeiT-small. For object detection and instance segmentation tasks, we use MMDetection [42] with a training pipeline aligned with SR-STE [10]. We employ SpRe to all N:M weights except for 1$\times$ 1 kernels wherein the notion of spatial sparsity is deemed inapplicable. Our experiments are implemented with PyTorch [43] and run on NVIDIA Tesla A100 GPUs.
4.2 Image Classification
CIFAR-10. We first evaluate the efficacy of SpRe for sparsifying ResNet-32 on the CIFAR-10 dataset, which includes 50,000 training images and 10,000 validation images within 10 classes. Table I shows that SpRe can be well leveraged to enhance the performance of all N:M methods. For instance, the Top-1 classification accuracy of SR-STE is improved by 0.2%, 0.3%, 0.9%, and 1.3% for 2:4, 1:4, 1:8, and 1:16 patterns, respectively. Given introducing no extra inference, the effectiveness of SpRe for compensating the limited variation on spatial sparsity of existing methods is obvious.
ImageNet-1K. For the large-scale ImageNet-1K dataset that contains over 1.2 million images for training and 50,000 images for validation in 1,000 categories, we first present the quantitative results for sparsifying ResNet [15] with depths of 18 in Table II and 50 in Table III. Encouragingly, SpRe substantially enlarges the performance of N:M methods over a wide range of sparse patterns. Without extra inference burden introduced, $0.2\%$ and $0.5\%$ Top-1 accuracy improvements are gained by equipping ASP [12] with SpRe when sparsifying ResNet-18 at 2:4 and 1:4 patterns. The advantage of SpRe becomes more evident with an increase in the sparsity rate, where classic N:M sparsity methods fail to allocate enough weights for processing the important visual points due to a constant spatial sparsity. Upon training 1:16 sparse ResNet-50, SpRe exhibits a remarkable enhancement in the Top-1 accuracy of SR-STE [10] and LBC [13] by 1.2% and 1.1%, respectively. The improvement is attributed to our weight preservation at crucial visual locations. Therefore, the efficacy of SpRe on the large-scale challenging tasks is also evident.
Furthermore, we examine the generalization capability of SpRe for sparsifying MobileNet-V1 [38] and MobileNet-V2, two lightweight networks incorporating depth-wise convolution design and posing more challenges for compression. Table IV and Table V respectively show the quantitative results. It can be seen that all N:M methods suffer significantly more performance drops for sparsifying lightweight networks. Nevertheless, SpRe demonstrates a distinct trend whereby the performance improvements remain consistent and increase as the sparsity level increases, irrespective of the specific N:M methods employed. For instance, SpRe is able to enhance the Top-1 accuracy of ASP by $0.3\%$, $2.5\%$, $4.5\%$and $7.6\%$ when sparsifying MobileNet-V1 at 2:4, 1:4, 1:18, 1:16 sparse patterns, respectively. Similar phenomenon can be drawn when it comes to MobileNet-V2. These results well highlight the efficacy of SpRe in advancing existing N:M sparsity methods for sparsifying lightweight networks.
TABLE IV: Results for sparsifying MobileNet-V1 at different N:M patterns on ImageNet-1K.
Model Method N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc
MobileNet-V1 Baseline - 70.9 - 70.9 - 70.9 - 70.9
MobileNet-V1 ASP [12] 2:4 70.2 1:4 63.9 1:8 61.4 1:16 50.4
MobileNet-V1 w. SpRe 2:4 70.5(+0.3) 1:4 66.6(+2.5) 1:8 65.9(+4.5) 1:16 58.0(+7.6)
MobileNet-V1 SR-STE [10] 2:4 70.4 1:4 63.2 1:8 52.1 1:16 25.4
MobileNet-V1 w. SpRe 2:4 70.8(+0.4) 1:4 64.9(+1.7) 1:8 54.8(+2.7) 1:16 34.2(+8.8)
TABLE V: Results for sparsifying MobileNet-V2 at different N:M patterns on ImageNet-1K.
Model Method N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc N:M Top-1 Acc
MobileNet-V2 Baseline - 72.5 - 72.5 - 72.5 - 72.5
MobileNet-V2 ASP [12] 2:4 71.0 1:4 67.4 1:8 64.0 1:16 58.1
MobileNet-V2 w. SpRe 2:4 71.3(+0.3) 1:4 67.9(+0.5) 1:8 65.2(+1.2) 1:16 60.4(+2.3)
MobileNet-V2 SR-STE [10] 2:4 69.5 1:4 63.2 1:8 53.3 1:16 24.2
MobileNet-V2 w. SpRe 2:4 69.9(+0.4) 1:4 64.3(+1.1) 1:8 55.0(+1.7) 1:16 35.4(+11.2)
TABLE VI: Performance comparison between N:M and unstructured sparsity methods for sparsifying ResNet-50 on ImageNet-1K.
Method Sparsity Top-1 Acc Structured
ResNet50 0 77.3 -
DNW [44] 90 74.0
RigL [19] 90 73.0
GMP [21] 90 73.9
STR [45] 91 74.0
GraNet [14] 90 74.2
SR-STE [10] 88(1:8) 73.8
w. SpRe 88(1:8) 74.7(+0.9)
LBC [13] 88(1:8) 74.0
w. SpRe 88(1:8) 74.8(+0.8)
DNW [44] 95 68.3
RigL [19] 95 70.0
GMP [21] 95 70.6
STR [45] 95 70.4
GraNet [14] 95 72.3
SR-STE [10] 94(1:16) 71.5
w. SpRe 94(1:16) 72.7(+1.2)
LBC [13] 94(1:16) 71.8
w. SpRe 94(1:16) 72.9(+1.1)
Additionally, we provide the performance comparison between unstructured sparsity and N:M sparsity methods in Table VI. The results reveal that N:M sparsity methods struggle to maintain accuracy on par with advanced unstructured sparsity techniques at comparable levels of sparsity. This disparity in performance can be attributed to the fact that unstructured sparsity, as we previously discussed, sustains adequate processing of crucial visual points at high levels of sparsity. In contrast, N:M sparsity fails to effectively handle this aspect due to its constant spatial sparsity and hence lags behind in achieving comparable outcomes. Fortunately, by allocating re-parameterizable weights that follow the spatial sparsity distribution of unstructured sparsity, SpRe effectively elevates the accuracy of N:M sparsity methods to the level of state-of-the-art unstructured sparsity methods without incurring any extra inference overhead. For instance, SR-STE surpasses the recent unstructured sparsity method GraNet [14] by 0.4% Top-1 accuracy with the aid of SpRe (72.7% for SR-STE boosted by SpRe and 72.3% for GraNet at similar sparsity rates). Given the distinct advantage of N:M sparsity for practical acceleration on the N:M sparse tensor core, the significance of SpRe in bridging the performance gap between N:M sparsity and unstructured sparsity is apparent.
TABLE VII: Results for object detection on COCO.
Model Method N:M  mAP
F-RCNN Baseline - 37.4
F-RCNN SR-STE 2:4 38.2
F-RCNN w. SpRe 2:4 38.4(+0.2)
F-RCNN SR-STE 1:4 37.1
F-RCNN w. SpRe 1:4 37.3(+0.2)
F-RCNN SR-STE 1:16 35.8
F-RCNN w. SpRe 1:16 36.3(+0.5)
4.3 Object Detection and Instance Segmentation
Beyond fundamental image classification benchmarks, we exploit the generalization ability of SpRe on the object detection and instance segmentation tasks of COCO benchmark [37]. Table VII compares our proposed SpRe to SR-STE for training N:M sparse Faster-RCNN [40]. Notably, SpRe yields robust performance improvement of $0.2$, $0.2$ and $0.5$ mAP at 2:4, 1:4, and 1:16 sparse patterns, respectively. Similar trends can be observed from Table VIII when sparsifying Mask-RCNN [41] for the instance segmentation task. These results well substantiate the robustness and effectiveness of SpRe on downstream computer vision tasks.
TABLE VIII: Results for instance segmentation on COCO.
Model Method N:M Box mAP Mask mAP
M-RCNN Baseline - 38.2 34.7
M-RCNN SR-STE 2:4 39.0 35.3
M-RCNN w. SpRe 2:4 39.2(+0.2) 35.7(+0.4)
M-RCNN SR-STE 1:4 37.4 33.5
M-RCNN w. SpRe 1:4 37.7(+0.3) 33.9(+0.4)
M-RCNN SR-STE 1:16 35.4 31.8
M-RCNN w. SpRe 1:16 36.1(+0.7) 32.5(+0.7)
4.4 Performance Analysis
Extra Branch. We first present the performance analysis of SpRe by investigating two variants on the derivation of the extra branch in Eq. (8). The experiments incorporate 2:4, 1:4, and 1:16 patterns to sparsify ResNet-50 on ImageNet-1K, with SR-STE [10] serving as the comparative baseline. Examining the details, we first perform ablation that maintains the same mask with the main N:M sparse weights as $\mathbf{B}^{S}=\mathbf{B}$. This variation carries out more preserved weights in the extra branch, despite the spatial sparsity mirroring that of the vanilla N:M sparsity. As depicted in Fig 5, more weights in the extra branch (referred to as Same) even bring less performance improvement. We attribute this phenomenon to limited variability in spatial sparsity as discussed in Sec. 3.2. Besides, we consider another variation that allocates parameters at spatial locations where unstructured sparsity exhibits less spatial sparsity than N:M sparsity (referred to as Inverse). As can be seen, the inverse allocation for the weights in the extra branch brings even negative performance, which further demonstrates our point that a correct spatial sparsity variation is the key to the performance retention of sparse networks.
Additional Overhead. We further elucidate the overhead implicated in Floating Point Operations (FLOPs) and parameters concomitant with the introduction of SpRe. Table IX enumerates the quantitative outcomes from employing SR-STE for the sparsification of ResNet-50 on ImageNet-1K. The deployment of SpRe does, indeed, incur a modicum of additional overhead during the training phase, albeit minimal. However, during the inference (test) stage, the quantity of parameters and FLOPs perseveres consistently, enabling efficacious deployment and acceleration founded on the Sparse Tensor Core [12]. Concurrently, juxtaposed with the results presented in Table IX, SpRe achieves a palpable enhancement in performance with the introduction of merely a negligible training overhead, affirming its undoubted practical utility.
Figure 5: Performance comparison for using different strategies to generate the extra branch. Experiments are based on using SR-STE to sparsify ResNet-50 on ImageNet-1k.
TABLE IX: FLOPs and parameters comparison for pruning ResNet-50 on ImageNet-1K at training and testing stages.
Method N:M FLOPs FLOPs Params Params
(Train) (Test) (Train) (Test)
ResNet-50 - 1$\times$ 1$\times$ 1$\times$ 1$\times$
SR-STE 2:4 0.83$\times$ 0.5$\times$ 1$\times$ 0.5$\times$
w. SpRe 2:4 1.02$\times$ 0.5$\times$ 1.24$\times$ 0.5$\times$
SR-STE 1:4 0.74$\times$ 0.25$\times$ 1$\times$ 0.25$\times$
w. SpRe 1:4 0.80$\times$ 0.25$\times$ 1.14$\times$ 0.25$\times$
SR-STE 1:8 0.71$\times$ 0.13$\times$ 1$\times$ 0.13$\times$
w. SpRe 1:8 0.73$\times$ 0.13$\times$ 1.09$\times$ 0.13$\times$
SR-STE 1:16 0.69$\times$ 0.06$\times$ 1$\times$ 0.06$\times$
w. SpRe 1:16 0.70$\times$ 0.06$\times$ 1.05$\times$ 0.06$\times$
5 Limitation
We further discuss unexplored limitations, which will be our future focus. SpRe is particularly presented for N:M sparsity in convolutional neural networks upon our observations of spatial sparsity variability. Although not the focus of this current work, it would be interesting for future work to examine similar discoveries to drive further enhancement for N:M sparsity in networks with different typologies, e.g., Vision Transformers (ViTs). Besides, Though improving the performance, SpRe also brings some extra burdens for N:M sparsity in the training time. A promising direction in the future is to develop more efficient ways to mitigate the lack in spatial sparsity variability of N:M sparsity.
6 Conclusion
In this work, we have presented Spatial Re-parameterization, an effective and easy-to-use method for N:M sparsity. By introducing a re-parameterizable branch that follows the spatial sparsity distribution of unstructured sparsity, SpRe is able to reimburse the spatial sparsity variability during the training time of N:M sparsity, which we examine to be the core for performance retention. Our proposed SpRe is benchmarked on several computer vision benchmarks, consistently delivering enhanced performance for representative N:M methods without extra inference burden. Notably, SpRe brings the performance of N:M sparsity methods to a comparable level with unstructured sparsity methods for the first time. Hopefully, this work shields a convincing trail to dive into the intrinsic property of N:M sparsity.
Acknowledgement
This work was supported by National Key R&D Program of China (No.2022ZD0118202), the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of China (No. U21B2037, No. U22B2051, No. 62176222, No. 62176223, No. 62176226, No. 62072386, No. 62072387, No. 62072389, No. 62002305 and No. 62272401), and the Natural Science Foundation of Fujian Province of China (No.2021J01002, No.2022J06001).